OpenStudy (anonymous):

plz help diffrentiate 2{cot(x^2)}^1/2

hartnn (hartnn):

first tell what is differentiation of x^1/2? and have u studied chain rule?

OpenStudy (anonymous):

1/2(x)^1/2

hartnn (hartnn):

right,so differentiation of cot (x^2)^1/2 will be?.....before applying chain rule..

OpenStudy (anonymous):

1/2(cotx2)^1/2

hartnn (hartnn):

right,now if u know chain rule,what will get multiplied to this??

OpenStudy (anonymous):

-sin(x^2)

OpenStudy (anonymous):

\[-2x (\csc^2 x^2) / \sqrt{\cot x^2}\] please check if I'm right.

OpenStudy (anonymous):

2x afterwards

hartnn (hartnn):

how sin (x^2) ?? what is d/dx of cot x ?? the 2x part afterwards is correct....

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i was doing that mistake only

OpenStudy (anonymous):

[-\csc^2 x\]

hartnn (hartnn):

according to standard formula its - cosec ^2 x

OpenStudy (anonymous):

ok thanks now i try it

hartnn (hartnn):

post the final answer u get,so that we can verify....

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

-2xcosec^(x)^2/{cot(x^2)}^1/2

hartnn (hartnn):

yup u are correct :)

OpenStudy (anonymous):

i got this answer aim was right but in ncert answer is -2\[\sqrt{2} x/\sin(x)^2\sqrt{\sin2x^2}\]